### v2

Abstract
In this paper, we study a class of real-valued mean-field backward stochastic differential equations (BSDEs) with generator of quadratic growth in the control variable and the mean-field term. Under this assumption, together with a bounded terminal condition, we establish existence and uniqueness of solutions. Our approach departs from classical fixed-point arguments and instead combines Malliavin calculus with refined BMO and stability estimates. The result bridges the gap between the quadratic BSDE results of cheridito2017bse and hao2022mean. Moreover, motivated by the structure of the mean-field exponential utility maximization problem introduced in our paper, we extend our framework to generators satisfying a weaker quadratic condition on the generator. This relaxation is designed to accommodate the additional mean-field terms that arise in our utility maximization setting and that fall outside the scope of previous quadratic assumptions. Within this more general regime, we establish existence and uniqueness of solutions under a smallness condition on the terminal random variable. We then apply this extended theory to solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility to a fully coupled quadratic mean-field setting.
Key words: mean-field BSDEs; quadratic BSDEs; BMO martingales; Malliavin calculus; exponential utility maximization.
MSC: 60H30; 60H07; 60H10.
1 Introduction
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a filtered probability space where $W$ is a $d$-dimensional Brownian motion and $\{\mathcal{F}_{t}\}_{t\geq 0}$ is the Brownian filtration. Consider the following mean-field BSDE:
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},$ (1.1)
[…]
The origin of linear BSDEs can be traced back to their role as adjoint equations in stochastic control problems, first studied by bismut1973conjugate. The seminal paper by pardoux1990adapted established the existence and uniqueness of adapted solutions $(Y,Z)$ for
$Y_{t}=\xi+\int_{t}^{T}F(s,Y_{s},Z_{s})\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$ (1.2)
under the conditions that the generator $F$ is Lipschitz continuous and the terminal condition $\xi$ is square integrable. Since then, there has been a growing interest in BSDEs with applications in various fields such as stochastic control, mathematical finance, and partial differential equations. For a comprehensive presentation of the theory, see el1997backward, CohenStocal2015 or ZhangBSDEbook2017. Numerous studies have aimed to improve existence and uniqueness results by relaxing the Lipschitz condition on the generator $F$ or the conditions on the terminal condition $\xi$.
BSDEs with generators exhibiting quadratic growth in the variable $z$ have attracted significant attention. These equations arise, for example, in stochastic linear-quadratic control with random coefficients bismut1976linearquad, in utility optimization problems in an incomplete market hu2005utility. Kobylanski:2000cy provided the seminal existence and uniqueness result for real-valued quadratic BSDEs with bounded terminal values, where $F$ has quadratic growth with respect to its $Z$ component. briand2013simple have continued to explore this area, demonstrating the existence of unique solutions to quadratic BSDEs with bounded terminal conditions using purely probabilistic methods. Investigations by briand2006bsde extend the previous results to unbounded terminal conditions. Some studies the multidimensional cases: see hu2016multi; tang2023MultiMFBSDE; luo2019triangle; tevzadze2008solvability.
Mean-field stochastic differential equations, also known as McKean-Vlasov equations, describe stochastic systems whose evolution is influenced by both the microscopic state and the macroscopic distribution of particles. These equations date back to kac1956foundations’s work on the Boltzmann and Vlasov equations in the 1950s. Mean-field game theory began with the groundbreaking contributions of lasry2007mfgame. Building on this foundation, Buckdahn2009MFlimit introduced mean-field BSDEs, derived from approximations involving weakly interacting $N$ particle systems, and investigated them using probabilistic methods. This has led to a growing interest in mean-field BSDEs and FBSDEs. For instance, Buckdahn2009MFBSDE studied the existence and uniqueness of solutions with Lipschitz continuous generators and square integrable terminal conditions. More recently, hao2022mean explored the solvability of solutions for one-dimensional mean-field BSDEs with quadratic growth, while tang2023MultiMFBSDE established the existence and uniqueness of solutions for multidimensional BSDEs with generators depending on the expectation of both variables. Li2018MFBSDEcontinuous examined mean-field BSDEs whose coefficients depend not only on the solution $(Y,Z)$ but also on the law of $Y$.
pardoux1990adapted interpreted the key structural feature of BSDEs as a nonlinear conditional expectation, highlighting their close link to Malliavin calculus. In fact, el1997backward and Buckdahn2009MFBSDE state that, in a Clark–Ocone type formula, the martingale density component $Z$ of the solution pair of a BSDE with a classical globally Lipschitz generator on a Gaussian basis is the Malliavin derivative of the other component $Y$ (see Proposition 5.3). The result has been extending our understanding of quadratic BSDEs: for example, briand2013simple provided a straightforward approach to the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus to overcome difficulties with quadratic generators, and Imkeller2010path proposed a numerical scheme for BSDEs with generators of quadratic growth.
Continuous time portfolio optimization under the utility-maximization framework introduced by merton1969portfolio is a foundational topic in mathematical finance. Merton analyzed the problem in a complete market where the risky asset follows the Black-Scholes dynamics and no constraints are imposed on admissible trading strategies. A major development in the early 2000s was the emergence of a connection between this class of optimization problems and the theory of BSDEs, which has since become a central approach in the field. The core idea of the BSDE methodology is to construct a stochastic process, dependent on the investment strategy, that coincides with the investor’s utility of terminal wealth. In this context, elrouge2000utility used BSDEs to compute the value function and optimal strategy for exponential utility, assuming trading strategies are restricted to a convex cone. Building on this framework, hu2005utility developed a BSDE approach to utility maximization in incomplete markets and extended the methodology to include logarithmic and power utility functions, allowing for general strategy constraints within a closed set.
Our methods are inspired by the works of tevzadze2008solvability and briand2013simple. The former studied the existence and uniqueness of solutions for multidimensional quadratic BSDEs, utilizing BMO martingale theory. The latter provided a straightforward approach to establishing the existence and uniqueness of quadratic BSDEs with bounded terminal conditions using Malliavin calculus. Building on these foundations, we bridge the gap between the approaches of cheridito2017bse and hao2022mean. Specifically, hao2022mean established the existence and uniqueness by applying a contraction mapping with the Reverse Hölder Inequality (RHI) and iterating backward in time when the generator is quadratic in $Z$ component, while cheridito2017bse used the Krasnoselskii fixed-point theorem on a small time interval, iterating backward to obtain a global solution. Our contribution extends the results of cheridito2017bse by demonstrating in Proposition 4.9 that the function $f^{1}$ satisfies the conditions outlined in hao2022mean, thereby proving the existence and uniqueness of a solution for quadratic mean-field BSDEs, which is quadratic both in $Z$ and the corresponding mean-field component $\tilde{Z}$.
Our main contributions are as follows. First, we establish the existence and uniqueness of a solution when the generator exhibits separately quadratic growth in both $Z$ and the mean-field term $\tilde{Z}$ with bounded terminal condition $\xi$. Our approach builds upon the theory of BMO martingales and Malliavin calculus to effectively address the analytical challenges arising from the quadratic structure. A key novelty of this work lies in the inclusion of a quadratic mean-field component $\tilde{Z}$, which distinguishes it from existing results in the literature. Moreover, our method is notable for its simplicity, as it avoids the use of fixed-point arguments commonly employed in previous studies, such as juan2024MFBSDE and tevzadze2008solvability.
The rest of the paper is organized as follows. In Section 2, we present the mathematical framework and state the main result. Section 3 derives an almost sure upper bound for the solution pair $(Y,Z)$ of Lipschitz mean-field BSDEs, assuming the terminal condition admits a uniformly bounded Malliavin derivative, as shown in Proposition 3.1. In the main Section 4, we provide a priori estimates and establish a powerful stability result for mean-field BSDEs with separately quadratic growth. In Section 5, we utilize the bounds obtained in Section 3, which are independent of the Lipschitz constants in $z$ and $\tilde{z}$, together with the stability result and a density argument, to prove the existence and uniqueness of solutions to our quadratic mean-field BSDEs. Finally, in Section 6, we extend the existence and uniqueness results for mean-field quadratic BSDE with separately quadratic driver to a fully coupled quadratic driver with small terminal conditions, and show how the corresponding BSDE characterizes a utility maximization problem with a particular class of constraint sets.
2 Preliminaries and statement of main results
2.1 Preliminaries
[…]
  • $L^{p}$: all real-valued random variables $X$ s.t. $\|X\|_{L^{p}}:=\left({\mathbb{E}}^{\mathbb{P}}\big[|X|^{p}\big]\right)^{\frac{1}{p}}<\infty$,
  • $L^{\infty}$: all essentially bounded random variables $X$ s.t. $\|X\|_{L^{\infty}}:=\operatorname*{ess\,sup}_{\omega\in\Omega}|X(\omega)|<\infty$,
  • $\mathbb{S}^{p}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
    $\left\|Y\right\|_{\mathbb{S}^{p}}:=\left(\mathbb{E}\left[\sup_{0\leq r\leq T}\left|Y_{r}\right|^{p}\right]\right)^{\frac{1}{p}}<\infty,$
  • $\mathbb{S}^{\infty}$: the set of real valued $\mathcal{F}$-adapted continuous processes $(Y_{t})_{0\leq t\leq T}$ satisfying
    $\left\|Y\right\|_{\mathbb{S}^{\infty}}:=\sup_{0\leq r\leq T}\left\|Y_{r}\right\|_{L^{\infty}}=\sup_{\omega\in\Omega}\sup_{0\leq r\leq T}\left|Y_{r}(\omega)\right|<\infty,$
  • $\mathbb{H}^{p}$: the set of predictable $\mathbb{R}^{d}$-valued process $Z$ satisfying $\left\|Z\right\|_{\mathbb{H}^{p}}:=\left(\mathbb{E}\left[\int_{0}^{T}\left|Z_{r}\right|^{2}\,dr\right]^{\frac{p}{2}}\right)^{\frac{1}{p}}<\infty,$
  • $\mathbb{H}^{2}_{BMO}$: a subset of $\mathbb{H}^{2}$ with $Z$ satisfying $Z\cdot W:=\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ is a Bounded Mean Oscillation (BMO) martingale, that is, there exists a non-negative constant $C$ such that, for each stopping time $\tau\leq T$,
    $\mathbb{E}\left[\int_{\tau}^{T}\left|Z_{r}\right|^{2}\,dr\Big|\mathcal{F}_{\tau}\right]\leq C^{2}.$
    The $\left\|Z\right\|_{\mathbb{H}^{2}_{BMO}}:=\left\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\right\|_{BMO}$ is defined as the smallest non-negative constant $C$ for which the above inequality is satisfied.
  • For a closed subset $C$ of $\mathbb{R}^{n}$ and $a\in\mathbb{R}^{n}$, the distance between $a$ and the set $C$ is defined by
    $\text{dist}(a,C)=\min_{b\in C}\left|a-b\right|.$
    The set $\Pi_{C}(a)$ consists of all points $b\in C$ at which this minimum is attained, and we refer to this as the projection set.
When a different underlying measure other than $\mathbb{P}$ is involved, we indicate it explicitly in the notation, for example by writing $\mathbb{E}$$^{\mathbb{Q}}$$\cdot$] and $\|\cdot\|_{BMO(\mathbb{Q})}$ for expectations and BMO norms taken with respect to a probability measure $\mathbb{Q}$. Here are some inequalities from BMO martingales theory, all the proofs can be found in kazamaki2006BMO.
Lemma 2.1 (Reverse Hölder Inequality (RHI)).
Let $M$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$. Then the Doléans–Dade exponential $\mathcal{E}(M)$ satisfies the reverse Hölder inequality: there exists $q^{*}>1$ such that for all $q\in(1,q^{*})$ and all stopping times $\tau\leq T$,
$\mathbb{E}\!\left[\left.\left(\frac{\mathcal{E}(M)_{T}}{\mathcal{E}(M)_{\tau}}\right)^{\!q}\;\right|\mathcal{F}_{\tau}\right]\;\leq\;C_{q}^{*},$ (2.3)
where the constants $q^{*}$ and $C_{q}^{*}$ depend explicitly on the $\mathrm{BMO}$ norm $\|M\|_{\mathrm{BMO}(\mathbb{P})}$:
$q^{*}\;:=\;\phi^{-1}\!\Big(\|M\|_{\mathrm{BMO}}\Big),\qquad\phi(x)\;:=\;\Big(1+\tfrac{1}{x^{2}}\log\frac{2x-1}{2x-2}\Big)^{1/2}-1,$
and for $1<q<q^{*}$,
$C_{q}^{*}\;:=\;2\Bigg(1-\frac{2q-2}{2q-1}\,\exp\!\Big\{q^{2}\|M\|_{\mathrm{BMO}}^{2}+2\|M\|_{\mathrm{BMO}}\Big\}\Bigg)^{-1}.$
In particular, $q^{*}$ decreases and $C_{q}^{*}$ increases with respect to $\|M\|_{\mathrm{BMO}}$.
Lemma 2.2.
For $K>0$. There exist constants $c_{1}(K),c_{2}(K)>0$, depending only on $K$, such that the following holds. Let $N$ be a one-dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ with $\|N\|_{\mathrm{BMO}(\mathbb{P})}\leq K$, and define the equivalent measure $\tilde{\mathbb{P}}$ by
$\frac{d\tilde{\mathbb{P}}}{d\mathbb{P}}\;=\;\mathcal{E}(N)_{T}.$
For any continuous $\mathrm{BMO}$ martingale $M$ under $\mathbb{P}$, set
$\tilde{M}\;:=\;M-\langle M,N\rangle,$
[…]
$c_{1}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}\;\leq\;\|\tilde{M}\|_{\mathrm{BMO}(\tilde{\mathbb{P}})}\;\leq\;c_{2}(K)\,\|M\|_{\mathrm{BMO}(\mathbb{P})}.$ (2.4)
Definition 2.3 (sliceability).
Let $M$ be a one–dimensional continuous $\mathrm{BMO}$ martingale under $\mathbb{P}$ on $[0,T]$. A random partition of $[0,T]$ is a finite sequence of stopping times $0=\tau_{0}\leq\tau_{1}\leq\cdots\leq\tau_{m}=T$. For $\delta>0$, the index of sliceability of $M$ is the function
$N_{M}:(0,\infty)\longrightarrow\mathbb{N}\cup\{\infty\}$
defined as follows: $N_{M}(\delta)$ is the smallest integer $m$ such that there exists a random partition $(t_{k})_{k=0}^{m}$ of $[0,T]$ with
$\big\|M^{(k)}\big\|_{\mathrm{BMO}}\leq\delta,\qquad 1\leq k\leq m,$ (2.5)
where the kk–th increment $M^{(k)}$ is the continuous martingale
$M^{(k)}_{t}:=M_{t\wedge\tau_{k}}-M_{t\wedge\tau_{k-1}},\qquad 0\leq t\leq T.$
If no such $m$ exists, we set $N_{M}(\delta)=\infty$. We say that $M$ is $\delta$–sliceable if $N_{M}(\delta)<\infty$, and simply sliceable if $M$ is $\delta$–sliceable for every $\delta>0$.
2.2 Statement of the main theorem
In this paper, we consider the following general mean-field BSDE
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T].$ (2.6)
The main result of this paper establishes the existence and uniqueness of solutions to the mean-field BSDE under a bounded terminal condition. To this end, we introduce the following assumptions. Let $K_{0},K_{z},K_{\tilde{z}}>0$ be fixed constants.
  1. $(\rm H_{0})$}
    The random variable $\xi$ is bounded almost surely by $K_{0}>0$, and likewise, the deterministic supremum $\sup_{t\in[0,T]}|g(t,0,0)|$ is also bounded by the same constant $K_{0}>0$.
  2. $(\rm H_{q})$}
    There exist constants $K_{z}>0$ and $K_{\tilde{z}}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
Example 2.4.
The following $g$ satisfies the condition $(\mathrm{H}_{q})$:
[…]
where $l$ is Lipschitz in (z,z~)($\tilde{z}$), and $q_{1}$ and $q_{2}$ are quadratic polynomials in $z$ and $\tilde{z}$, respectively. Note that this example does not satisfy the existence and uniqueness assumption in previous literature, such as hao2022mean.
Theorem 2.5.
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
Remark 2.6.
In the statement and proof of our main theorem we may, without loss of generality, we can assume that
$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
This is not an additional structural condition on the original generator, but merely a convenient normalisation. Given any generator $g$ satisfying Assumptions $(\rm H_{0})$$(\rm H_{q})$}, we may introduce the shifted generator
[…]
and consider the mean-field BSDE driven by $\hat{g}$. For any solution $(Y,Z)$ of the original BSDE with generator $g$, the process
$U_{t}:=Y_{t}-\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds,\qquad 0\leq t\leq T,$
together with the same $Z$, solves the BSDE with generator $\hat{g}$. Conversely, given a solution $(U,Z)$ to the BSDE with generator $\hat{g}$, the process
$Y_{t}:=U_{t}+\int_{t}^{T}\widetilde{{\mathbb{E}}}\big[g(s,0,\tilde{Z}_{s})\big]\,ds$
solves the original BSDE. Thus there is a one-to-one correspondence in terms of existence and uniqueness. From this point on, we therefore work directly with the shifted generator, still denoted by $g$, and silently assume $g(t,0,\tilde{z})=0$ for all $(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}$.
3 Terminal conditions with bounded Malliavin derivative
This section is devoted to deriving tractable upper bounds for the solution $(Y,Z)$ of the BSDE (2.6) under the assumption that the driver $g$ is Lipschitz continuous and the terminal condition $\xi$ is Malliavin differentiable. These bounds under the stronger condition become particularly useful in the proof of Lemma 5.1, we will show that the general case, where the driver exhibits quadratic growth and where $\xi$ is bounded, can be reduced to this setting through suitable approximation techniques. Therefore, we initially work under stronger regularity assumptions to facilitate this reduction. For this purpose, we suppose all along this section that $g$ satisfies the following classical Lipschitz condition and condition on $\xi$ will be introduced in Propisiton 3.1 below.
  • $({\rm H_{lip}})$
    There exist a constant $L_{z}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
The existence and uniqueness of a solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ to the BSDE (2.6) under assumptions $(\rm H_{0})$$({\rm H_{lip}})$ is a well-established result, as shown in Theorem 3.1 of Buckdahn2009MFBSDE. In the following proposition, we demonstrate that the process $Y$, which lies in $\mathbb{S}^{\infty}$, is bounded by a constant that is independent of the Lipschitz constant $L_{z}$. More remarkably, we show that the same independence holds for the process $Z$, provided the terminal condition $\xi$ admits uniformly bounded Malliavin derivatives.
Proposition 3.1.
If $(\rm H_{0})$} and $({\rm H_{lip}})$ hold, then the unique solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ of the Lipschitz mean-field BSDE satisfies the following properties.
  1. 1.
    $Y\in\mathbb{S}^{\infty}$ with the upper bound:
    $\|Y\|_{\mathbb{S}^{\infty}}\;\leq\;\left\|\xi\right\|_{L^{\infty}}.$ (3.7)
  2. 2.
    If $\xi$ is Malliavin differentiable and satisfies $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}=\sup_{0\leq t\leq T}\left\|D_{t}\xi\right\|_{L^{\infty}}<\infty$, then $Z\in\mathbb{S}^{\infty}$, and its norm is also bounded independently of $L_{z}$:
    $\left\|Z\right\|_{\mathbb{S}^{\infty}}\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}},$ (3.8)
Proof.
Unlike standard Lipschitz BSDE, the comparison theorem cannot be applied here (see Theorem 3.2 in Buckdahn2009MFBSDE). Observe that (3.7) can be derived following the line of proof of Proposition 4.4. In this Lipschitz context, the proof is simpler since no BMO martingale arguments are necessary.
[…]
$D_{\theta}Y_{t}=D_{\theta}\xi+\int_{t}^{T}\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]\!\cdot D_{\theta}Z_{s}\,ds-\int_{t}^{T}D_{\theta}Z_{s}\!\cdot dW_{s}.$
Under assumption $({\rm H_{lip}})$, we have $\left|\partial_{z}g\right|\leq L_{z}$. Therefore, the Novikov condition is satisfied and $\widetilde{{\mathbb{E}}}\!\left[\partial_{z}g(s,Z_{s},\widetilde{Z}_{s})\right]$ is measurable to $mathcal{F_{s}}$, and we may apply the Girsanov transform to obtain the following representation with new Brownian motion $W^{\mathbb{Q}}$ under the new probability measure $\mathbb{Q}$:
$D_{\theta}Y_{t}=D_{\theta}\xi-\int_{t}^{T}D_{\theta}Z_{s}\!\cdot dW_{s}^{\mathbb{Q}},\qquad 0\leq\theta\leq t\leq T.$
Taking the conditional expectation under $\mathbb{Q}$, we deduce
$\left|D_{\theta}Y_{t}\right|\leq\left\|D\xi\right\|_{\mathbb{S}^{\infty}}.$
Finally, as $(Z_{t})_{0\leq t\leq T}$ is a version of $(D_{\theta}Y_{t})_{0\leq t\leq T}$ and due to the arbitrariness of $t$, we establish 3.8. ∎
4 A priori estimates and stability for BSDEs
In this section, we will work under the assumptions $(\rm H_{0})$} and $(\rm H_{q})$} to establish the stability of the mean-field BSDE (2.6), as stated in Theorem 4.6. The strategy of the proof is as follows. We first derive a priori upper bounds on the $\mathbb{S}^{\infty}$ norm of $Y$ and the BMO norm of $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ for any solution $(Y,Z)$ to the BSDE (2.6). These estimates, together with fundamental properties of BMO martingales, lead to the stability result. We begin with a useful estimate that will play an important role in the analysis.
Remark 4.1.
Under assumptions $(\rm H_{0})$} and $(\rm H_{q})$}, we have the following estimate:
[…]
for any $(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{d}\times\mathbb{R}^{d}$. Taking expectations with respect to $\tilde{\mathbb{P}}$ yields:
$\tilde{\mathbb{E}}\left|g(t,Z_{t},\tilde{Z}_{t})\right|\leq K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}+\frac{3K_{z}}{2}\left|Z_{t}\right|^{2}+\frac{3K_{\tilde{z}}}{2}\tilde{\mathbb{E}}\left|\tilde{Z}_{t}\right|^{2}.$ (4.9)
We begin by noting that if the solution $Y\in\mathbb{S}^{\infty}$, then for any solution pair $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ to the quadratic mean-field BSDE, the stochastic integral $\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}$ defines a BMO martingale. Furthermore, the BMO norm of this martingale can be explicitly bounded in terms of the uniform norm $\left\|Y\right\|_{\mathbb{S}^{\infty}}$ and certain constants in the assumptions.
Proposition 4.2.
Suppose that (H0) and (Hq) hold. Let $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$ be a solution to the mean-field BSDE (2.6). Then the stochastic process $\left(\int_{0}^{t}Z_{s}\!\cdot dW_{s}\right)_{0\leq t\leq T}$ defines a BMO martingale. Moreover, its BMO norm satisfies the following bound:
$\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}^{2}\leq(1\vee T)\big(4K_{0}+K_{z}+K_{\tilde{z}}\big)\,\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\left(\frac{K_{\tilde{z}}}{K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\right).$ (4.10)
Proof.
The argument follows from arguments in Proposition 2.1 of briand2008stobsdes with modified $C^{2}$ function
$\varphi(x):=\frac{e^{3K_{z}|x|}-1-6cK_{z}|x|}{(3K_{z})^{2}},\qquad c\in\mathbb{R},x\in\mathbb{R}.$
In Step 1 we express the $\mathrm{BMO}$ norm of $Z$ in terms of its $\mathbb{H}^{2}$ norm; in Step 2 we derive an explicit $\mathbb{H}^{2}$ bound in terms of the model coefficients and $\|Y\|_{\mathbb{S}^{\infty}}$.
Step 1 (Estimate of $\|Z\|_{\mathbb{H}^{2}_{\mathrm{BMO}}}$ in terms of $\|Z\|_{\mathbb{H}^{2}}$). Fix $c=\tfrac{1}{2}$ then $\varphi(0)=0$ and, for all $x\in\mathbb{R}$,
$\varphi(x)\geq 0,\qquad\varphi^{\prime}(x)=\frac{e^{3K_{z}|x|}-1}{3K_{z}}\operatorname{sgn}(x),\qquad\varphi^{\prime\prime}(x)=e^{3K_{z}|x|},\qquad\varphi^{\prime\prime}(x)-3K_{z}\,|\varphi^{\prime}(x)|\equiv 1.$
Moreover, for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$ we have the deterministic bound
$0\leq|\varphi^{\prime}(x)|\leq C_{Y}:=\frac{1}{3K_{z}}\,e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}.$ (4.11)
Let $\tau$ be any stopping time and apply Itô’s formula to $\varphi(Y)$ and using the growth estimate 4.9 for $g$ together with the symmetry $\widetilde{\mathbb{E}}|\tilde{Z}_{s}|^{2}=\mathbb{E}|Z_{s}|^{2}$ yields
$\varphi(Y_{\tau})$ $=\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(-\varphi^{\prime}(Y_{s})\,\widetilde{\mathbb{E}}[g(s,Z_{s},\tilde{Z}_{s})]-\tfrac{1}{2}\,\varphi^{\prime\prime}(Y_{s})|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}$
$\leq\varphi(\xi)+\!\int_{\tau}^{T}\!\Big(\frac{3K_{z}}{2}|\varphi^{\prime}(Y_{s})|-\tfrac{1}{2}\varphi^{\prime\prime}(Y_{s})\Big)|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds$
$\leq\varphi(\xi)-\tfrac{1}{2}\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds+\!\int_{\tau}^{T}\!|\varphi^{\prime}(Y_{s})|\Big(K+\tfrac{3}{2}K_{\tilde{z}}\,\mathbb{E}|Z_{s}|^{2}\Big)ds-\!\int_{\tau}^{T}\!\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s},$
where $K:=K_{0}+\frac{1}{2}K_{z}+\frac{1}{2}K_{\tilde{z}}$. Taking $\mathbb{E}[\cdot\mid\mathcal{F}_{\tau}]$ and using (4.11) together with $|\varphi(x)|\leq C_{Y}|x|$ (since $\varphi(0)=0$ and $\varphi^{\prime}$ is bounded on the range of $Y$), we get
$\tfrac{1}{2}\,\mathbb{E}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]$ $\leq\mathbb{E}\!\big[\varphi(\xi)-\varphi(Y_{\tau})\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}K(T-\tau)+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds$
$\leq C_{Y}\,\mathbb{E}\!\big[|\xi|\,\big|\,\mathcal{F}_{\tau}\big]+C_{Y}KT+\tfrac{3}{2}K_{\tilde{z}}C_{Y}\!\!\int_{0}^{T}\!\mathbb{E}|Z_{s}|^{2}ds.$
we thus obtain
$\tfrac{1}{2}\,{\mathbb{E}}\!\Big[\!\int_{\tau}^{T}\!|Z_{s}|^{2}ds\,\Big|\,\mathcal{F}_{\tau}\Big]\leq C_{Y}\left(\|\xi\|_{L^{\infty}}+K\,T+\tfrac{3}{2}K_{\tilde{z}}\!\!\int_{0}^{T}\!\!{\mathbb{E}}|Z_{s}|^{2}ds\right).$ (4.12)
Since $\tau$ is arbitrary, (4.12) shows that $\big\|\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\big\|_{\mathrm{BMO}}^{2}$ is controlled by $\|Z\|_{\mathbb{H}^{2}}^{2}$ and model parameters.
Step 2 (An explicit $\mathbb{H}^{2}$ bound). Write $\alpha:=K_{\tilde{z}}/K_{z}>0$. We may, without loss of generality, assume $\alpha$ satisfies
$0<\alpha\;<\;\min\!\Big\{\tfrac{1}{2},\ \frac{e^{3K_{z}}-2}{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}\Big\}.$
by enlarging $K_{z}$ or $K_{\tilde{z}}$. Define positive constants
$c_{1}:=\frac{K_{\tilde{z}}}{2K_{z}}e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+\frac{1}{2},\qquad c_{2}:=\tfrac{1}{2}e^{3K_{z}}.$
Note that, by definition of $\alpha$, $c_{1}$ can also be expressed as $c_{1}=\tfrac{1}{2}\big(\alpha e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}+1\big)$. By the choice of $\alpha$, it is easy to verify $c_{1}<c_{2}$. Pick any $c\in(c_{1},c_{2})$ and consider the same $\varphi$ at in Step 1 with a different $c$ value, so that $\varphi(0)=0$ and, for all $x$,
$|\varphi^{\prime}(x)|=\frac{e^{3K_{z}|x|}-2c}{3K_{z}}>\frac{e^{3K_{z}|x|}-2c_{2}}{3K_{z}}\geq 0.$
We also have identity $\varphi^{\prime\prime}(\cdot)-3K_{z}\left|\varphi^{\prime}(\cdot)\right|=2c$. Applying Itô to $\varphi(Y)$ on $[0,T]$ and arguing as in Step 1 with $\tau=0$ gives
$\varphi(Y_{0})$ $\leq\varphi(\xi)-\int_{0}^{T}c\,|Z_{s}|^{2}ds+\frac{3K_{\tilde{z}}}{2}\int_{0}^{T}|\varphi^{\prime}(Y_{s})|\,\mathbb{E}|Z_{s}|^{2}ds+K\int_{0}^{T}|\varphi^{\prime}(Y_{s})|ds-\int_{0}^{T}\varphi^{\prime}(Y_{s})Z_{s}\cdot dW_{s}.$
Taking expectations and using $|\varphi^{\prime}(x)|=\frac{e^{3K_{z}x}-2c}{3K_{z}}\leq\frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}=C_{Y}$ for $|x|\leq\|Y\|_{\mathbb{S}^{\infty}}$,
$\varphi(Y_{0})\leq\mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T-c\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]+\frac{3K_{\tilde{z}}}{2}\,C_{Y}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right].$
By the definition of $c_{1}$ and the fact that $c>c_{1}=\frac{3K_{\tilde{z}}}{2}C_{Y}+\tfrac{1}{2}$ we have $\frac{3K_{\tilde{z}}}{2}C_{Y}<c-\tfrac{1}{2}$. Hence
$\frac{1}{2}\,\mathbb{E}\!\left[\int_{0}^{T}|Z_{s}|^{2}ds\right]\ \leq\ \mathbb{E}\!\left[\varphi(\xi)\right]+KC_{Y}T\ \leq\ \frac{e^{3K_{z}\|Y\|_{\mathbb{S}^{\infty}}}}{3K_{z}}\,\big(\left\|\xi\right\|_{L^{\infty}}+KT\big).$ (4.13)
Finally, plugging (4.13) into (4.12) and taking the essential supremum over $\tau$ yields an explicit $\mathrm{BMO}$ bound for $\int_{0}^{\cdot}Z_{s}\cdot dW_{s}$ in terms of $\|Y\|_{\mathbb{S}^{\infty}}$ and the model parameters. This completes the proof. ∎
Next, suppose the solution $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$. Then, for any solution pair $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the quadratic mean-field BSDE, we derive the following proposition showing that $Y\in\mathbb{S}^{\infty}$ which is independent of the BMO norm $\left\|\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\right\|_{\mathrm{BMO}}$. We introduce a simple but useful lemma which will be used later on when we work under equivalent probability measures.
Lemma 4.3.
Let $\mathbb{P}$ and $\mathbb{Q}$ be equivalent probability measure on $(\Omega,\mathcal{F})$. For any random variable $X$,
$\|X\|_{L^{\infty}(\mathbb{P})}=\|X\|_{L^{\infty}(\mathbb{Q})}.$
In particular, $X$ is bounded $\mathbb{P}$-a.s. iff it is bounded $\mathbb{Q}$-a.s., with the same bound.
Proof.
Set $K:=\|X\|_{L^{\infty}(\mathbb{P})}$. For every $\varepsilon>0$, $\mathbb{P}(|X|>K+\varepsilon)=0$. Since $\mathbb{Q}\ll\mathbb{P}$, also $\mathbb{Q}(|X|>K+\varepsilon)=0$, so $\|X\|_{L^{\infty}(\mathbb{Q})}\leq K$. By symmetry ($\mathbb{P}\ll\mathbb{Q}$), the reverse inequality holds. ∎
Proposition 4.4.
Suppose Assumptions $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then for any solution $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ of BSDE (2.6), we have $Y\in\mathbb{S}^{\infty}$ and the following estimate holds:
$\left\|Y\right\|_{\mathbb{S}^{\infty}}\leq\left\|\xi\right\|_{L^{\infty}}.$ (4.14)
Proof.
The proof relies on a linearization argument combined with the BMO property. Using Remark 2.6, without the loss of generality, we assume
$g(t,0,\tilde{z})=0,\qquad\forall\,(t,\tilde{z})\in[0,T]\times\mathbb{R}^{d}.$
Then, the BSDE (2.6) can be rewritten as
$Y_{t}=\xi+\int_{t}^{T}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot Z_{s}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad t\in[0,T],$ (4.15)
where the process $(L_{s})_{0\leq s\leq T}$ is defined by
$L_{s}:=\frac{g(s,Z_{s},\tilde{Z}_{s})-g(s,0,\tilde{Z}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}}\qquad 0\leq s\leq T,$
Since $g(s,0,\tilde{z})=0$ by construction, the numerator is simply $g(s,Z_{s},\tilde{Z}_{s})$. Moreover, Assumption $(\rm H_{q})$} yields
$\Big|\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big|\leq\int_{\tilde{\Omega}}\left|L_{s}\right|\,d\tilde{\mathbb{P}}\leq K_{z}\big(1+\left|Z_{s}\right|\big),\qquad 0\leq s\leq T.$
Since $Z\in\mathbb{H}^{2}_{\mathrm{BMO}}$ by assumption, $\big(\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\cdot dW_{s}\big)_{t\in[0,T]}$ is a BMO martingale, and its stochastic exponential $\mathcal{E}_{t}$ satisfies Kazamaki’s condition. Define the probability measure $\mathbb{P}^{L}$ by $d\mathbb{P}^{L}:=\mathcal{E}_{T}\,d\mathbb{P}$. By the above BMO argument, $(\mathcal{E}_{t})$ is a uniformly integrable martingale, so $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$. Girsanov’s theorem then implies that the process
$W_{t}^{L}:=W_{t}+\int_{0}^{t}\Big(\int_{\tilde{\Omega}}L_{s}\,d\tilde{\mathbb{P}}\Big)\,ds,\qquad t\in[0,T],$
is a dd–dimensional Brownian motion under $\mathbb{P}^{L}$, thus we obtain
$Y_{t}=\xi-\int_{t}^{T}Z_{s}\!\cdot dW_{s}^{L}.$
Hence,
$Y_{t}={\mathbb{E}}^{\mathbb{P}^{L}}\big[\xi\,\big|\,\mathcal{F}_{t}\big],\qquad 0\leq t\leq T.$
Since $\mathbb{P}^{L}$ is equivalent to $\mathbb{P}$, the same bound holds $\mathbb{P}$–almost surely by Lemma 4.3, and taking the essential supremum over $t\in[0,T]$ completes the proof. ∎
Then we have a direct result from the previous two propositions.
Corollary 4.5.
Let $(\rm H_{0})$} and $(\rm H_{q})$} hold. Then, for any solution pair $(Y,Z)$ to the quadratic BSDE (2.6), if either
$(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}\quad\text{or}\quad(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}},$
it follows that $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ and
$\left\|Y\right\|_{\mathbb{S}^{\infty}}+\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\leq C,$ (4.16)
where $C$ is a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$.
Proof.
We only sketch how the two estimates interact, as the details follow directly from Propositions 4.2 and 4.4. If $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}$, then Proposition 4.2 yields
$Z\in\mathbb{H}^{2}_{\mathrm{BMO}}\quad\text{and}\quad\Big\|\textstyle\int_{0}^{\cdot}Z_{s}\!\cdot dW_{s}\Big\|_{\mathrm{BMO}}\ \text{is bounded in terms of }\|Y\|_{\mathbb{S}^{\infty}}.$
Since we are now in the setting of Proposition 4.4, we obtain an $\mathbb{S}^{\infty}$ estimate for $Y$ which depends only on $\left\|\xi\right\|_{L^{\infty}}$. Combining these two facts shows that both $\|Y\|_{\mathbb{S}^{\infty}}$ and the BMO norm of $\int Z\,dW$ are bounded by some constant $C$ depending only on the model parameters ($K_{0},K_{z},K_{\tilde{z}},T$). The other case $(Y,Z)\in\mathbb{S}^{2}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ is the same. ∎
Theorem 4.6.
For $i=1,2$, let $(Y^{i},Z^{i})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$ solve
$Y_{t}^{i}=\xi^{i}+\int_{t}^{T}\!\int_{\tilde{\Omega}}g(s,Z_{s}^{i},\tilde{Z}_{s}^{i})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{i}\cdot dW_{s},\qquad t\in[0,T],$
with $\xi^{1},\xi^{2},g$ satisfying $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. Then there exist $p^{*}>1$ and, for every $p>p^{*}$, a constant $C_{p}>0$ such that
$\|Y^{1}-Y^{2}\|_{\mathbb{S}^{2p}}+\|Z^{1}-Z^{2}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{1}-\xi^{2}\|_{L^{2p}},$
where $p^{*}>1$ and $C_{p}$ depend only on $T$, $K_{z}$, $K_{\tilde{z}}$, $K_{0}$, and do not depend on $\|Z^{i}\|_{\mathbb{S}^{\infty}}$.
Proof.
Let $\Delta Y:=Y^{1}-Y^{2}$, $\Delta Z:=Z^{1}-Z^{2}$, $\Delta\xi:=\xi^{1}-\xi^{2}$ and
$\Delta g(s,Z_{s},\tilde{Z}_{s}):=g(s,Z_{s}^{1},\tilde{Z}_{s}^{1})-g(s,Z_{s}^{2},\tilde{Z}_{s}^{2}),\qquad s\in[0,T].$
Then we obtain
$\Delta Y_{t}=\Delta\xi+\int_{t}^{T}\!\int_{\tilde{\Omega}}\!\Delta g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}\Delta Z_{s}\cdot dW_{s},\qquad t\in[0,T].$ (4.17)
Preliminary step (linearization, RHI and sliceability). As in the proof of Proposition 4.4, we use a linearization argument: there exist processes $L_{s}$ and $\tilde{L}_{s}$ on the product space such that, for all $s\in[0,T]$,
$L_{s}:=\frac{g(s,Z^{1}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})}{|\Delta Z_{s}|^{2}}\,\Delta Z_{s}\,\mathbf{1}_{\{\Delta Z_{s}\neq 0\}},\qquad\tilde{L}_{s}:=\frac{g(s,Z^{2}_{s},\tilde{Z}^{1}_{s})-g(s,Z^{2}_{s},\tilde{Z}^{2}_{s})}{|\Delta\tilde{Z}_{s}|^{2}}\,\Delta\tilde{Z}_{s}\,\mathbf{1}_{\{\Delta\tilde{Z}_{s}\neq 0\}}$
with estimates
$|L_{s}|\leq K_{z}\big(1+|Z_{s}^{1}|+|Z_{s}^{2}|\big),\qquad|\tilde{L}_{s}|\leq K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big),$ (4.18)
and $\Delta g(s,Z_{s},\tilde{Z}_{s})=L_{s}\Delta Z_{s}+\tilde{L}_{s}\,\Delta\tilde{Z}_{s}$.
According to Proposition 4.2, we have $Z^{1},Z^{2}\in\mathbb{H}^{2}_{\mathrm{BMO}}$, so in particular $\big(\int_{0}^{t}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\big)_{t\in[0,T]}$ is a $\mathrm{BMO}$ martingale by (4.18) and symmetry. Moreover, by Corollary 4.5, there exists a constant $C>0$, depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$, such that
$\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{1}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}+\Big\|\int_{0}^{\cdot}Z^{2}_{s}\cdot dW_{s}\Big\|_{\mathrm{BMO}}\;\leq\;C.$ (4.19)
Hence $\Lambda:=\mathcal{E}(\int_{0}^{\cdot}\widetilde{\mathbb{E}}[L_{s}]\cdot dW_{s})$ defines an equivalent probability measure $\mathbb{P}^{L}$ by Girsanov’s theorem with Brownian motion $W^{L}_{t}:=W_{t}+\int_{0}^{t}\widetilde{\mathbb{E}}[L_{s}]\,ds$. In addition, by Lemma 2.2, the same processes admit $\mathrm{BMO}$ norms under the equivalent measure $\mathbb{P}^{L}$, up to multiplicative constants depending only on $C$. Under $\mathbb{P}^{L}$ we can rewrite (4.17) as
$\Delta Y_{u}=\Delta\xi+\int_{u}^{T}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{T}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in[0,T].$ (4.20)
The Doléans–Dade exponential $\Lambda=\mathcal{E}\!\left(\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\right)$ satisfies a reverse Hölder inequality by Lemma 2.1: there exists $q^{*}>1$ such that, for all $1\leq q\leq q^{*}$,
$\bar{{\mathbb{E}}}\!\left[\left(\Lambda_{T}\right)^{q}\,\big|\,\bar{\mathcal{F}}_{t}\right]\ \leq\ C_{q}^{*}\,\left(\Lambda_{t}\right)^{q},\qquad 0\leq t\leq T,$
where
$q^{*}$ $:=\phi\!\left(\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right),\qquad\phi(x):=\left(1+\frac{1}{x^{2}}\log\frac{2x-1}{2x-2}\right)^{\!1/2}-1,$
$C_{q}^{*}$ $:=2\left(1-\frac{2q-1}{2q-2}\,\exp\!\left\{q^{2}\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+2\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}\right\}\right)^{-1}\!.$
By boundedness of $\Big\|\int_{0}^{\cdot}\tilde{\mathbb{E}}[L_{s}]\cdot dW_{s}\Big\|_{\mathrm{BMO}}$, the constants $q^{*},C_{q}^{*}$ depend only on $(T,K_{z},K_{\tilde{z}},K_{0})$. Finally, set
$M_{t}:=\int_{0}^{t}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s},\qquad t\in[0,T].$
Then $M\in\mathcal{H}^{\infty}(\mathbb{P})$, and by schachermayer1996institut it is sliceable. Furthermore, since $Z^{1},Z^{2}\in\mathbb{S}^{\infty}$, we can (by (jackson2022triangle, Remark 2.8)) construct a deterministic partition $0=t_{0}<t_{1}<\dots<t_{N}=T$ with arbitrarily small mesh size bounded by a prescribed $\varepsilon>0$.
$\mathbb{S}^{p}$–control of $\Delta Y$. Fix $\varepsilon$, $k\in\{0,\dots,N-1\}$ and write $I_{k}=[t_{k},t_{k+1}]$. By subtracting (4.20) at $u=t_{k+1}$ from that at $u\in I_{k}$, we obtain the local BSDE
$\Delta Y_{u}=\Delta Y_{t_{k+1}}+\int_{u}^{t_{k+1}}\widetilde{\mathbb{E}}[\tilde{L}_{s}\,\Delta\tilde{Z}_{s}]\,ds-\int_{u}^{t_{k+1}}\Delta Z_{s}\cdot dW^{L}_{s},\qquad u\in I_{k}.$
Using Cauchy-Schwarz on the copy space, the growth bound (4.18), and then Fefferman’s inequality (Theorem 2.5 in kazamaki2006BMO) on $I_{k}$, we obtain
$\int_{u}^{t_{k+1}}\tilde{\mathbb{E}}|\tilde{L}_{s}||\Delta\tilde{Z}_{s}|ds$ $\leq\mathbb{E}\int_{0}^{T}(\tilde{\mathbb{E}}\big|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)\big|^{2})^{1/2}(\tilde{\mathbb{E}}|\Delta\tilde{Z}_{s}|^{2})^{1/2}ds$ (4.21)
$\leq\sqrt{2}\left\|\int_{0}^{\cdot}(\tilde{\mathbb{E}}|K_{\tilde{z}}\big(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|\big)|^{2})^{1/2}dW_{s}\right\|_{BMO}\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds$
$\leq\varepsilon\int_{u}^{t_{k+1}}(\mathbb{E}|\Delta Z_{s}|^{2})^{1/2}ds\qquad u\in I_{k}.$
Taking $\mathbb{E}^{L}[\cdot\,|\,\mathcal{F}_{u}]$ in the local BSDE yields
$|\Delta Y_{u}|$ $\leq\mathbb{E}^{L}[|\Delta Y_{t_{k+1}}|\,|\,\mathcal{F}_{u}]+\varepsilon\int_{u}^{t_{k+1}}\big(\mathbb{E}|\Delta Z_{s}|^{2}\big)^{1/2}ds,\qquad u\in I_{k}.$
Iterating this estimate over the intervals $I_{k+1},I_{k+2},\dots,I_{N-1}$ and using $\Delta Y_{T}=\Delta\xi$, we obtain
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]+\varepsilon\sqrt{T}\,\left(\int_{u}^{T}\mathbb{E}|\Delta Z_{s}|^{2}\,ds\right)^{1/2},\qquad u\in[0,T].$
By BMO energy inequality (kazamaki2006BMO, Sec. 2.1) and Corollary 4.5, the last term is bounded by a constant depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ since:
$\int_{u}^{T}\Big({\mathbb{E}}|Z_{s}^{1}|^{2}+{\mathbb{E}}|Z_{s}^{2}|^{2}\Big)\,ds\;$ $\leq\;\Big\|\int_{0}^{\cdot}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}+\Big\|\int_{0}^{\cdot}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2}.$
Therefore, letting $\varepsilon\to 0$ yields
$|\Delta Y_{u}|\leq\mathbb{E}^{L}[|\Delta\xi|\,|\,\mathcal{F}_{u}]=(\Lambda_{u})^{-1}\,\mathbb{E}[\Lambda_{T}|\Delta\xi|\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$ (4.22)
Now fix $q\in(1,q^{*})$ and let $p$ be its conjugate exponent. By the reverse Hölder inequality and conditional Hölder’s inequality,
$|\Delta Y_{u}|^{p}$ $\leq(\Lambda_{u})^{-p}\,\mathbb{E}[(\Lambda_{T})^{q}\,|\,\mathcal{F}_{u}]^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}]$
$\leq(C_{q}^{*})^{p/q}\,\mathbb{E}[|\Delta\xi|^{p}\,|\,\mathcal{F}_{u}],\qquad u\in[0,T].$
By Doob’s maximal inequality,
$\mathbb{E}\Big(\sup_{u\in[0,T]}|\Delta Y_{u}|^{p}\Big)\ \leq\ C_{Y}(p)\,\mathbb{E}|\Delta\xi|^{p},\qquad p\geq p^{*},$ (4.23)
where $p^{*}$ is the conjugate exponent of $q^{*}$ and $C_{Y}(p):=\Big(\frac{p}{p-1}\Big)^{p}(C_{q}^{*})^{p/q}$. They both depending only on $(T,K_{z},K_{\tilde{z}},K_{0})$ as we discussed in the Preliminary step.
$\mathbb{H}^{p}$–control of $\Delta Z$. Fix $p>p^{*}$, and applying Itô’s formula to $|\Delta Y|^{2}$ on $[0,T]$. we obtain
$|\Delta Y_{0}|^{2}+\!\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds=|\Delta\xi|^{2}+2\int_{0}^{T}\Delta Y_{s}\,\widetilde{\mathbb{E}}[\Delta g(s,Z_{s},\tilde{Z}_{s})]\,ds-2\int_{0}^{T}\Delta Y_{s}\,Z_{s}\cdot dW_{s}$
Using the same Fefferman/Cauchy–Schwarz argument as above and $\|\Delta Y\|_{\mathbb{S}^{\infty}}<\infty$, we obtain
$2\int_{0}^{T}\Delta Y_{s}\,\widetilde{{\mathbb{E}}}\!\big[\Delta g(s,Z_{s},\tilde{Z}_{s})\big]\,ds$ $\leq 2K_{z}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)\,|\Delta Y_{s}|\,|\Delta Z_{s}|\,ds$
$\quad+2K_{\tilde{z}}\int_{0}^{T}|\Delta Y_{s}|\,\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds$
$\leq\sup_{t\in[0,T]}|\Delta Y_{t}|^{2}\Big(2K_{z}^{2}\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds\Big)+\frac{1}{2}\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds$
$\quad+2K_{\tilde{z}}\|\Delta Y\|_{\mathbb{S}^{\infty}}\int_{0}^{T}\widetilde{{\mathbb{E}}}\!\Big[(1+|\tilde{Z}_{s}^{1}|+|\tilde{Z}_{s}^{2}|)\,|\Delta\tilde{Z}_{s}|\Big]\,ds.$
The last term vanishes in the limit $\varepsilon\to 0$ by the same slicing argument as in (4.21). Next, let $K:=\int_{0}^{T}(1+|Z_{s}^{1}|+|Z_{s}^{2}|)^{2}\,ds$. Then
${\mathbb{E}}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\,K^{p/2}\Big]\leq\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[K^{p}\big]\Big)^{1/2}.$ (4.24)
If $\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}$ is a BMO martingale ($i=1,2$), the BMO energy inequality (see (kazamaki2006BMO, Sec. 2.1)) gives
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|Z_{s}^{i}|^{2}ds\Big)^{p}\Big]\leq p!\,\Big\|\int_{0}^{\cdot}Z_{s}^{i}\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}.$
Therefore,
${\mathbb{E}}[K^{p}]\leq 3^{p-1}\Big((3T)^{p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{1}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}+3^{p}p!\,\Big\|\int_{0}^{t}Z_{s}^{2}\,\cdot dW_{s}\Big\|_{\mathrm{BMO}}^{2p}\Big)=:C_{\mathrm{energy}}(p).$ (4.25)
Lastly, with the help of BDG inequality and Cauchy–Schwarz, we have
${\mathbb{E}}\Big|\int_{0}^{T}(\Delta Y_{s})\Delta Z_{s}\cdot dW_{s}\Big|^{p/2}$ $\leq C_{\mathrm{BDG}}\,{\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Y_{s}|^{2}|\Delta Z_{s}|^{2}\,ds\Big)^{p/4}\Big]$
$\leq C_{\mathrm{BDG}}\,\Big({\mathbb{E}}\big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\big]\Big)^{1/2}\Big({\mathbb{E}}\big[(\int_{0}^{T}|\Delta Z_{s}|^{2}\,ds)^{p/2}\big]\Big)^{1/2}.$ (4.26)
Putting (4.24), (4.25) and (4) together, we obtain
${\mathbb{E}}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}{\mathbb{E}}[|\Delta\xi|^{p}]+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\Big({\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}]\Big)^{1/2}$
$\quad+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,{\mathbb{E}}[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}],$
From previous step (4.23), there exist constants $C_{Y}(p),C_{Y}(2p)$ such that
$\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{p}\Big]\leq C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big],\qquad\Big(\mathbb{E}\Big[\sup_{t\in[0,T]}|\Delta Y_{t}|^{2p}\Big]\Big)^{1/2}\leq\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Substituting these bounds into the previous inequality, we obtain for all $p\geq p^{*}$,
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq 2^{\frac{3p}{2}-1}\mathbb{E}\big[|\Delta\xi|^{p}\big]+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)\,\mathbb{E}\big[|\Delta\xi|^{p}\big]$
$\quad+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2}.$
Finally, since $\mathbb{E}|\Delta\xi|^{p}\leq\big(\mathbb{E}|\Delta\xi|^{2p}\big)^{1/2}$, we may absorb the first two terms into the last one and conclude that
$\mathbb{E}\Big[\Big(\int_{0}^{T}|\Delta Z_{s}|^{2}ds\Big)^{p/2}\Big]$ $\leq C_{Z}(p)\,\Big(\mathbb{E}\big[|\Delta\xi|^{2p}\big]\Big)^{1/2},$
where
$C_{Z}(p):=2^{\frac{3p}{2}-1}+2^{4p-4}(C_{\mathrm{BDG}})^{2}\,C_{Y}(p)+2^{\frac{3p}{2}-1}(2K_{z}^{2})^{p/2}\sqrt{C_{\mathrm{energy}}(p)}\,\sqrt{C_{Y}(2p)}.$
Combing the stability estimate (4.23) completes the proof. ∎
5 Proof: existence and uniqueness
In this section, we establish the main result for the mean-field BSDE (2.6) under Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$. The proof proceeds in two main steps. First, we prove the existence of a solution when the terminal condition $\xi$ is Malliavin differentiable with bounded derivative. This is achieved in Lemma 5.1 via a truncation and regularization argument. With the help of the estimates for Lipschitz BSDEs derived in Proposition 3.1, this allows us to apply the classical result from Buckdahn2009MFBSDE.
In the second step, we extend this existence result to the case of general bounded terminal conditions by approximating $\xi$ with a sequence of Malliavin differentiable random variables. For each approximating BSDE, existence is guaranteed by Lemma 5.1, and we invoke the stability result from Theorem 4.6 to show that the sequence of solutions is Cauchy in suitable norms. The limit of this sequence is then shown to solve the original mean-field BSDE (2.6). Uniqueness follows directly from the same stability result. The full argument is completed in the proof of the main theorem below.
Lemma 5.1.
Let $(\rm H_{0})$} and $(\rm H_{q})$} holds and suppose $\xi$ is Malliavin differentiable with bounded derivative $\left\|D\xi\right\|_{\mathbb{S}^{\infty}}<\infty$, then the mean-field BSDE admits a solution in $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}.$
Proof.
We define a sequence of driver functions $(\bar{g}_{k})_{k\in\mathbb{N}}$ as follows:
$\bar{g}_{k}(t,z,\tilde{z}):=g\left(t,\frac{\left|z\right|\wedge k}{\left|z\right|}z,\frac{\left|\tilde{z}\right|\wedge k}{\left|\tilde{z}\right|}\tilde{z}\right),\quad k\in\mathbb{N}.$
Since $g$ satisfies Assumption $(\rm H_{q})$}, it is straightforward to verify that the new driver $\bar{g}_{k}$ is Lipschitz, with
[…]
for all $k\in\mathbb{N}$ and $(t,z,z^{\prime},\tilde{z},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$. Thus, the corresponding BSDE with driver $\bar{g}_{k}$ satisfies $(\rm H_{0})$} and $(H_{\mathrm{lip}})$. Then, the mean-field BSDE
$Y_{t}^{k}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}\bar{g}_{k}(s,Z_{s}^{k},\tilde{Z}_{s}^{k})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{k}\cdot dW_{s},\quad 0\leq t\leq T,$ (5.27)
admits a unique solution $(Y^{k},Z^{k})\in\mathbb{S}^{2}\times\mathbb{H}^{2}$ (see Theorem 3.1 in Buckdahn2009MFBSDE). Moreover, Proposition 3.1 ensures that for each $k\in\mathbb{N}$, the pair $(Y^{k},Z^{k})$ belongs to $\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, and the process $Z^{k}$ satisfies a uniform bound independent of the Lipschitz constant of the driver with respect to $z$ and $\tilde{z}$. In particular, there exists a constant $C_{Z}>0$ such that $\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}\leq C_{Z}$ for all $k\in\mathbb{N}$. Since $(\tilde{Y}^{k},\tilde{Z}^{k})$ are independent copies of $(Y^{k},Z^{k})$, the same uniform bound holds, we can choose $k^{*}\in\mathbb{N}$ large enough so that
$k^{*}>\sup_{k\in\mathbb{N}}\left\|Z^{k}\right\|_{\mathbb{S}^{\infty}}.$
For such $k^{*}$, the truncation is inactive and
$\bar{g}_{k^{*}}(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}})=g(t,Z_{s}^{k^{*}},\tilde{Z}_{s}^{k^{*}}),\quad 0\leq s\leq T.$
Hence, the pair $(Y^{k^{*}},Z^{k^{*}})$ indeed solves the mean-field BSDE (2.6), which completes the proof. ∎
We are now in a position to state the main result of this section. For the reader’s convenience, we restate the main theorem below. To complete the proof of following theorem, it remains to extend the existence result from Malliavin differentiable terminal conditions to the general bounded case. Uniqueness then follows directly from the stability result established in Theorem 4.6.
Theorem 5.2.
Let Assumptions $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q})$ hold. Then, there exists a unique solution $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$ to the BSDE (2.6).
Proof.
We begin by approximating the bounded terminal condition $\xi$ with a sequence of Malliavin differentiable random variables of the form $\xi^{n}=\Phi^{n}(W_{t_{1}},\dots,W_{t_{n}})$, where each $\Phi^{n}$ is a bounded continuous function. Since $\xi$ is bounded, the sequence $(\xi^{n})_{n\in\mathbb{N}}$ can be chosen such that $\xi^{n}$ converges to $\xi$ in in any $L^{p}$ space for $p\geq 1$.
To ensure that the approximating BSDEs satisfy Assumption $(\mathrm{H}_{0})$, we truncate each $\Phi^{n}$ by setting $\Phi^{n}:=\Phi^{n}\wedge\left\|\xi\right\|_{L^{\infty}}$. Hence, each $\xi^{n}$ remains bounded, and $(\xi^{n},g)$ satisfies $(\mathrm{H}_{0})$ for all $n\in\mathbb{N}$. By Lemma 5.1, each approximating BSDE with terminal condition $\xi^{n}$ admits a solution $(Y^{n},Z^{n})\in\mathbb{S}^{\infty}\times\mathbb{S}^{\infty}$, satisfying
$Y_{t}^{n}=\xi^{n}+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s}^{n},\tilde{Z}_{s}^{n})\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}^{n}\cdot dW_{s}.$ (5.28)
Furthermore, the stability result in Theorem 4.6 implies that
$\|Y^{n}-Y^{m}\|_{\mathbb{S}^{2p}}+\|Z^{n}-Z^{m}\|_{\mathbb{H}^{p}}\ \leq\ C_{p}\;\|\xi^{n}-\xi^{m}\|_{L^{2p}}$
where $C_{p}$ depends only on $(T,K_{z},K_{\tilde{z}},K_{0})$, thus is independent of $n$ and $m$.
Pick $(\xi^{n})_{n\in\mathbb{N}}$ is Cauchy in $L^{2p}$, it follows that $(Y^{n},Z^{n})_{n\in\mathbb{N}}$ is also Cauchy in $\mathbb{S}^{2p}\times\mathbb{H}^{p}$. As each $(Y^{n},Z^{n})$ lies in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, we conclude the existence of a limit $(Y,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$. Passing to the limit in (5.28), we verify that the pair $(Y,Z)$ indeed solves the original mean-field BSDE (2.6), establishing existence. Uniqueness then follows directly from the stability result in Theorem 4.6, which concludes the proof. ∎
6 Application: Utility Maximization
In this section, we provide an extension of the foundational work of hu2005utility to the mean-field case. Specifically, we explore how a trader’s decision-making process evolves in a financial market influenced by collective behavior. Unlike previous studies that assume fixed constraints, we consider a dynamic setting where a trader’s strategy is shaped not only by individual preferences but also by the actions of others. This interaction leads to a more flexible and realistic model, where part of the trading constraints are given by market conditions rather than being imposed arbitrarily.
This section is structured as follows: In Section 6.1, we introduce the utility maximization problem and the associated mathematical framework. We then extend the given constraint set to incorporate a dependence on the stochastic process. It is important to note that, at this stage, we do not make any assumptions about this process beyond considering it as a general stochastic process. Next, we establish the connection between the utility maximization problem and the BSDE (6.30) with a thoughtfully chosen driver. We will subsequently demonstrate that the process defined earlier for the constraint set is precisely the solution $Z$ in the BSDE (6.30). With the complete framework in place, we offer a financial interpretation of the components, which is discussed in Remarks 6.4.
To connect this application to our earlier analysis about BSDE, we generalize the separately quadratic driver from BSDE (2.6) to a fully coupled quadratic driver in BSDE (6.32) and establish the existence and uniqueness of solutions in Proposition 6.7 when the terminal condition is small enough. To do this, we adopt a localization method in 6.6, then we extend it to the global case. Then, after we established the existence and uniqueness for BSDE (6.32), we then connect this back to the utility maximization problem by presenting the final result in Theorem 6.9. We first derive a key estimate in Lemma 6.3. This estimate is then used in Theorem 6.9 to verify that the driver of BSDE (6.30) satisfies conditions $(\rm H_{0})$} and $(\rm H_{q}^{\prime})$}, ensuring the well-posedness of BSDE (6.32). Leveraging these lemmas, we conclude with the main result in Theorem 6.9, which is supported by the proof of Theorem 7 in hu2005utility.
6.1 Financial market model setup
Consider a financial market composed of $d$ risky stocks. In this market, there are countably infinite number of homogeneous investors who trade dynamically between the risk-free bond and the risky assets. An investor, a representative of homogeneous investors, uses the following model to maximize the utility. The price process of the $i$-th stock, denoted by $S^{i}$, evolves according to the following stochastic differential equation:
$\frac{\mathrm{d}S_{t}^{i}}{S_{t}^{i}}=b_{t}^{i}\,\mathrm{d}t+\sigma_{t}^{i}\,\mathrm{d}W_{t},\quad i=1,\dots,d,$
where $W$ is $n$-dimensional Brownian motion, $b^{i}$ is a real-valued predictable, bounded stochastic process, representing the drift, and $\sigma^{i}$ is an $\mathbb{R}^{n}$-valued predictable bounded stochastic process, representing the volatility. The volatility matrix $\sigma_{t}=(\sigma_{t}^{1},\dots,\sigma_{t}^{d})^{\top}$ is assumed to have full rank $d$. And we assume that $\sigma_{t}\sigma_{t}^{\top}$ is uniformly elliptic. Thus market price of risk $\theta_{t}$, also referred to as the risk premium, is defined by:
$\theta_{t}:=\sigma_{t}^{\top}(\sigma_{t}\sigma_{t}^{\top})^{-1}b_{t},\quad t\in[0,T],$
where $\theta_{t}$ is an $\mathbb{R}^{n}$-valued uniformly bounded predictable process.
For $1\leq i\leq d$, let $\pi^{i}_{t}$ represent the amount of money invested in stock $i$ at time $t$, meaning the number of shares held in stock $i$ is $\frac{\pi^{i}_{t}}{S^{i}_{t}}$. A predictable process $\pi=(\pi_{t}^{1},\dots,\pi_{t}^{d})_{0\leq t\leq T}\in\mathbb{R}^{d}$, or simply $(\pi_{t})_{0\leq t\leq T}$, is called a trading strategy if $\int_{0}^{\cdot}\pi_{u}\frac{\mathop{}\!dS_{u}}{S_{u}}$ is well-defined, which holds if $\int_{0}^{T}\|\pi_{t}\sigma_{t}\|^{2}\mathop{}\!dt<\infty$ almost surely. The corresponding wealth process $X^{\pi}$, representing the investor’s wealth over time until $T$, is given by:
$X^{\pi}_{t}=x+\sum_{i=1}^{d}\int_{0}^{t}\frac{\pi^{i}_{u}}{S^{i}_{u}}\mathop{}\!dS^{i}_{u}=x+\int_{0}^{t}\pi_{u}\sigma_{u}(\mathop{}\!dW_{u}+\theta_{u}\mathop{}\!du),$
where $x\in\mathbb{R}$ is the initial capital.
At maturity $T$, the investor’s total wealth is determined by both their terminal wealth $X^{\pi}_{T}$ and an $\mathcal{F}_{T}$-measurable random liability $F$, where $F\geq 0$ represents a payment, and $F\leq 0$ represents income. The total wealth at maturity is thus $X^{\pi}_{T}-F$. The investor’s preferences are modeled by an exponential utility function applied to their terminal wealth $X^{\pi}_{T}-F$. The exponential utility function is given by:
$U(x)=-\exp(-\alpha x),\quad x\in\mathbb{R}.$
For the investors other than the representative, we assume the risk aversion parameter $\alpha>0$ is identical. The model for the risky assets is of the same type but with $b,\sigma,W$ being independent and identically distributed to the representative’s model. In addition, we assume the terminal liabilities are independent and have distribution $F$.
6.2 Constraint set and aggregated market signal
To ensure that no arbitrage opportunities exist, the definition of admissible trading strategies imposes constraints on the set of allowable positions. Below, we outline the constraint set in Definition 6.1, which accounts for market-wide influences. To incorporate the influence of market-wide behavior, for $t\in[0,T]$ and $\omega\in\Omega$, we define two stochastic processes that satisfy the necessary integrability conditions such that
$Z_{t}(\omega):=\hat{Z_{t}}(\omega)\sigma_{t}(\omega),$
where $Z_{t}(\omega)\in\mathbb{R}^{1\times n}$ and $\hat{Z_{t}}(\omega)\in\mathbb{R}^{1\times d}$. These two stochastic processes are essentially equivalent because the matrix $\sigma$ has full rank. Later on, in Remark 6.4, we will see $\hat{Z}_{t}\sigma_{t}+\frac{\theta_{t}}{\alpha}=Z_{t}+\frac{\theta_{t}}{\alpha}$ yields the optimal strategy at time $t\in[0,T]$ if no constants are imposed. Thus $Z$ can be interpreted as the trading strategy adjusted by $\theta/\alpha$. Later, we will show that this $Z$ will precisely correspond to the martingale density process of BSDE (6.30) as well.
We now present the following expressions for the constraint set under different settings. Specifically, we denote the constraint without market influence by $\hat{D}\subseteq\mathbb{R}^{d}$, where $\hat{D}$ is a closed set. When market influence is present, the strategy $\pi_{t}(\omega)$ is constrained to a new closed set.
The expectation of $\hat{Z}$ is referred to as the aggregated market signal, representing the expected collective adjusted trading strategy of all market participants. This signal plays a crucial role in defining the feasible trading region (constraint set), as described below.
Definition 6.1 (Mean-field adaptive constraint set).
Let $\hat{D}$ be a closed set in $\mathbb{R}^{d}$, and define the mean-field adaptive constraint set as:
$\hat{C}(\mathbb{E}[\hat{Z}_{t}]):=\hat{D}\odot L(\mathbb{E}[\hat{Z}_{t}]),\quad t\in[0,T],$ (6.29)
where $\odot$ denotes the Hadamard product (component-wise multiplication) and $L(\mathbb{E}[\hat{Z}_{t}])$ is a linear mapping depending on the value of $\mathbb{E}[\hat{Z}_{t}]$ for fixed $t\in[0,T]$, representing how the trader adjusts decisions based on the market’s behavior, defined by:
$L(\mathbb{E}[\hat{Z}_{t}]):=(\gamma\odot\mathbb{E}[\hat{Z}_{t}])+\eta,$
where $\gamma:=(\gamma_{1},\dots,\gamma_{d})$ is the market sensitivity vector and $\eta\in\mathbb{R}^{1\times d}$ is the market shift vector, with both being constant vectors.
With the help of the previous definitions, for a fixed $t\in[0,T]$, we start with a set $\hat{D}\subset\mathbb{R}^{d}$, which represents certain constraints, such as risk limits or position constraints. Upon incorporating information from the market, we could apply scaling and shifting to this set, which now reflects both the original constraints and the market’s influence. This results in a new set that is much more flexible and realistic, capturing the dynamic behavior of the market. Therefore, we denote this updated set as $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$, emphasizing its dependence on the aggregated market signal $\mathbb{E}[\hat{Z_{t}}$] at time $t\in[0,T]$.
Definition 6.2 (Admissible strategies with constraints $\hat{\mathcal{A}}$).
Let $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ be a closed set in $\mathbb{R}^{d}$ for fixed $t\in[0,T]$. The set of admissible trading strategies $\hat{\mathcal{A}}$ consists of all $\mathbb{R}^{d}$-valued predictable processes $\pi$ satisfying $\pi_{t}\sigma_{t}\in L^{2}[0,T]$, $\pi_{t}\in\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ a.s. $\forall t\in[0,T]$, and
[…]
forms a uniformly integrable family.
The investor’s objective is to choose an admissible self-financing trading strategy $\pi^{*}$ to maximize the expected utility of the net wealth at maturity $T$:
$V(0,x):=\sup_{\pi\in\hat{\mathcal{A}}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}\pi_{u}\frac{dS_{u}}{S_{u}}-F\right)\right)\right],$
where $V(0,\cdot)$ is called the value function at initial time 0. Solving this optimization problem requires selecting an admissible set from which we select the optimal trading strategy $\pi^{*}$. To make this notation more compact, we introduce the following notations. For $t\in[0,T],\omega\in\Omega$, define the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\subseteq\mathbb{R}^{n}$ and (volatility-scaled) trading strategy $p_{t}$ such that
$C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right):=\hat{C}(\mathbb{E}[\hat{Z}_{t}])\,\sigma_{t}(\omega)\quad\text{and}\quad p_{t}:=\pi_{t}\sigma_{t}.$
Note $Z_{t}=\hat{Z_{t}}\sigma_{t}$ and $\mathbb{E}[Z_{t}]=\mathbb{E}[\hat{Z_{t}}]\sigma_{t}$.
Lemma 6.3.
For every $(\omega,t)$, the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ is closed, and we have the following estimate:
$\max\{|c|:c\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\}\leq K_{1}\left(1+\mathbb{E}\left|Z_{t}\right|\right)\quad\text{a.s., for all }t\in[0,T],$
with some constant $K_{1}\geq 0$.
Proof.
Since the market sensitivity vector $\gamma$ and the market shift vector $\eta$ are both constant, the process $\sigma$ is uniformly bounded, and $\hat{D}$ is closed, the result follows. ∎
Note that the admissible trading strategy set $\mathcal{A}$, which is equivalent to $\hat{\mathcal{A}}$ ($\sigma$ has full rank), consists of all $\mathbb{R}^{n}$-valued predictable processes $p$ that satisfy $p\in L^{2}[0,T]$ and $p_{t}(\omega)\in C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ almost surely for all $t\in[0,T]$. Additionally, the family
[…]
is uniformly integrable. Finally, the dependency of the set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ on the process $(\hat{Z}_{t})_{0\leq t\leq T}$ is encoded by the dependency of the set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ on the process $(Z_{t})_{0\leq t\leq T}$. We keep the notation $X^{p}$ for the wealth process corresponding to investing strategy $p$.
With the help of these notations, we can rewrite the problem as
$\sup_{p\in\mathcal{A}}\mathbb{E}\left[-\exp\left(-\alpha\left(x+\int_{0}^{T}p_{t}(\mathop{}\!dW_{t}+\theta_{t}\mathop{}\!dt)-F\right)\right)\right]=\mathbb{E}\left[-\exp(-\alpha(X_{T}^{p^{*}}-F))\right]$
for the optimal control $p^{*}\in\mathcal{A}$. Let us introduce a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ as follows: let
$V^{(p)}(t,x):=-\exp(-\alpha(X_{t}^{p}-Y_{t})),\quad t\in[0,T],\,p\in\mathcal{A},$
where $(Y,Z)$ is a solution of the BSDE
$Y_{t}=F+\int_{t}^{T}f(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],.$ (6.30)
Here, the driver $f$ to be chosen such that $V^{(p)}(\cdot,x)$ is a supermartingale for all $p\in\mathcal{A}$, and there exists $p^{*}\in\mathcal{A}$ such that $V^{(p^{*})}(\cdot,x)$ is a true martingale. If we are able to choose such $f$, which will be shown later, we have
$\mathbb{E}V^{(p)}(T,x)\leq V^{(p)}(0,x)=V^{(p^{*})}(0,x)=\mathbb{E}V^{(p^{*})}(T,x).$
for all $p\in\mathcal{A}$ and the optimal control $p^{*}\in\mathcal{A}$.
Earlier, we described the process $(Z_{t})_{0\leq t\leq T}$ simply as a stochastic process. However, in this context, we now define it as the martingale density process $(Z_{t})_{0\leq t\leq T}$ for the BSDE in equation (6.30). Consequently, as the constraint set depends on $\mathbb{E}[Z_{t}]$, the driver should has dependence on $\mathbb{E}[Z_{t}]$ as well: see (6.31). To verify this, let us specify the driver $f$ for the construction of $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$. By Itô’s formula,
$V^{(p)}(t,x)=V^{(p)}(0,x)\tilde{A}^{p}\mathcal{E}_{t}\left(\int_{0}^{\cdot}\alpha(Z_{s}-p_{s})\cdot\mathop{}\!dW_{s}\right),$
where
$\tilde{A}^{p}=-\exp\left(\int_{0}^{t}\alpha v(s,p_{s},Z_{s})\mathop{}\!ds\right)$
and
$v(t,p_{t},Z_{t})=f(t,Z_{t},\mathbb{E}[Z_{t}])-p_{t}\theta_{t}+\frac{\alpha}{2}\left|p_{t}-Z_{t}\right|^{2}.$
Note that
$\mathcal{E}_{t}\left(\int_{0}^{\cdot}\alpha(Z_{s}-p_{s})\cdot\mathop{}\!dW_{s}\right)$
is a true martingale as $\alpha(Z-p)$ is square integrable. We require $\tilde{A}^{p}$ to be a non-increasing process such that $v(t,p_{t},Z_{t})\geq 0$ for all $p\in\mathcal{A}$, and there exists a $p^{*}\in\mathcal{A}$ such that $v(t,p^{*}_{t},Z_{t})=0$. Therefore, if we choose
$f(t,Z_{t},\mathbb{E}[Z_{t}])=-\frac{\alpha}{2}\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)+Z_{t}\theta_{t}+\frac{\left|\theta_{t}\right|^{2}}{2\alpha},$ (6.31)
then we have
$v(t,p_{t},Z_{t})$ $=f(t,Z_{t},\mathbb{E}[Z_{t}])+\frac{\alpha}{2}\left|p_{t}\right|^{2}+\frac{\alpha}{2}\left|Z_{t}\right|^{2}-\alpha p_{t}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$
$=f(t,Z_{t},\mathbb{E}[Z_{t}])+\frac{\alpha}{2}\left|p_{t}-\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)\right|^{2}-Z_{t}\theta_{t}-\frac{\left|\theta_{t}\right|^{2}}{2\alpha}$
$=\frac{\alpha}{2}\left|p_{t}-\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)\right|^{2}-\frac{\alpha}{2}\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)\geq 0.$
Thus, a family of processes $\left\{V^{(p)}(\cdot,x):p\in\mathcal{A}\right\}$ becomes a colletion of supermartingales where, for $p_{t}^{*}\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right)$, we have $v(\cdot,p^{*},Z)=0$ for every $t\in[0,T]$. Therefore, $V^{(p^{*})}(\cdot,x)$ becomes a true martingale.
We have now completed the construction, and it is clear how the utility maximization problem in our setting connects to a type of mean-field type BSDEs as in hibon2017MFquadratic. In hu2005utility, the optimization problem was solved under the assumption that the random endowment $F$ is bounded and that the trading strategy $p$ takes values in the admissible set $\mathcal{A}$. However, this set differs from ours because, in their case, the constraint set for trading strategies is simply closed in $\mathbb{R}^{d}$. In contrast, we introduce a particular constraint set where the trader’s behavior depends on the market (i.e., the behavior of other traders) in a mean-field sense. Specifically, we expand the driver to include the expectation of the control process $Z$. Thus, we define a constraint set that depends on time, reflecting the market’s influence at each moment. This approach introduces greater flexibility and a more realistic model. To establish the connection between the control process $Z$ in the BSDE (6.30) and the trading strategy $p\in\mathcal{A}$, we provide the following remarks, which bridge the mathematical framework with financial interpretations.
Remark 6.4.
If we take $\hat{D}$ to be the entire space, i.e., $\hat{D}=\mathbb{R}^{d}$, then the resulting constraint set $\hat{C}(\mathbb{E}[\hat{Z}_{t}])$ is again $\mathbb{R}^{d}$ as well. Consequently, we obtain $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)=\mathbb{R}^{n}$, implying the absence of constraints. In this case, the optimal trading strategy is given by
$p_{t}^{*}=Z_{t}+\frac{\theta_{t}}{\alpha},\quad\text{a.s., for all }t\in[0,T].$
Thus, the process $Z_{t}=p_{t}^{*}-\frac{\theta_{t}}{\alpha}$ can be interpreted as the unconstrained trading strategy adjusted by $\frac{\theta_{t}}{\alpha}$ at time $t\in[0,T]$. In general situation, the constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ depends on the average of the adjusted trading strategies among the market participants. The convergence is guaranteed by the homogeneity and exchangeability of market participants.
6.3 Existence and uniqueness for BSDEs with aggregated market signal
We are now ready to present the final result. We first borrow some ideas from tevzadze2008solvability, under small terminal value and lipschitz coefficient, the mean-field BSDE 6.32 admits a unique adapted solution. Next, in Proposition 6.7 we demonstrate that the driver in BSDE 6.30 guarantees the existence and uniqueness of solution. Finally, we complete this section by presenting the maximum expected utility value and corresponding trading strategy in in Theorem 6.9.
For the purpose of the final result, we introduce the BSDE (6.32) with the same formulation and the same $(\rm H_{0})$} condition in the main result but with a fully coupled quadratic growth condition and assume the following BSDE with driver $g$ and terminal condition $\xi$, satisfying the conditions specified below:
$Y_{t}=\xi+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\mathop{}\!d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\cdot\mathop{}\!dW_{s},\quad t\in[0,T].$ (6.32)
  • $(\rm H_{q}^{\prime})$}
    There exist constants $L_{z}>0$, and $K_{z}>0$ such that:
[…]
    for any $(t,z,\tilde{z},z^{\prime},\tilde{z}^{\prime})\in[0,T]\times[\mathbb{R}^{d}\times\mathbb{R}^{d}]^{2}$.
This is the fully coupled quadratic case, in contrast with the separately quadratic $(\rm H_{q})$} of the previous sections. We provide an explicit example here.
Example 6.5.
For simplicity, consider the one-dimensional case $d=1$ and define
$g(t,z,\tilde{z}):=z\,\tilde{z},\qquad(t,z,\tilde{z})\in[0,T]\times\mathbb{R}^{2}.$
This $g$ satisfies the fully coupled quadratic condition $(\rm H_{q}^{\prime})$}, but does not satisfy the separately quadratic condition $(\rm H_{q})$}.
We begin by establishing existence and uniqueness of solutions on a sufficiently small time interval in Lemma 6.6, where the required smallness conditions for the terminal value and time horizon are satisfied. By concatenating these local solutions, we extend the result to the global interval $[0,T]$ in Proposition 6.7.
First, we note that the terminal condition in BSDE 6.30 is $F$, representing a random liability. It is not realistic to assume $F$ is close to zero. Instead, we relax this requirement and require that $F$ is close to its mean. Accordingly, we rewrite the general BSDE 6.32 in terms of the centered random variable $\eta$:
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}\big[g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\big]\,d\tilde{\mathbb{P}}\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T,$ (6.33)
where
$U_{t}:=Y_{t}-g(t,0,0)-\mathbb{E}[\xi],\qquad\eta:=\xi-\mathbb{E}[\xi]$
are both well-defined because of $(\rm H_{0})$}.
Lemma 6.6.
Let $T<(64L_{z}^{2})^{-1}$ and assume that the driver $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$ and terminal condition $\eta$ be a bounded terminal condition such that
$\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
where
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}}.$
Then the BSDE 6.33 admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$, where
$\mathcal{B}_{\rho}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq\rho^{2}\right\}.$
Proof.
Step 1. We first establish existence and uniqueness for the auxiliary BSDE:
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\qquad 0\leq t\leq T.$
For each $(u,z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, with $\tilde{z}$ a copy of $z$, define $(U,Z)=\Gamma(u,z)$ to be the solution to
$U_{t}=\eta+\int_{t}^{T}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s}.$
Applying Itô’s formula to $|U_{t}|^{2}$ on $[t,T]$, we obtain
$|U_{t}|^{2}+\int_{t}^{T}|Z_{s}|^{2}ds=|\eta|^{2}+\int_{t}^{T}2U_{s}\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)\,d\tilde{\mathbb{P}}ds-2\int_{t}^{T}U_{s}\cdot Z_{s}\!\cdot dW_{s}.$
Taking the conditional expectation and applying $2ab\leq\frac{1}{2}a^{2}+2b^{2}$, we deduce
$|U_{t}|^{2}+\mathbb{E}\left[\int_{t}^{T}|Z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$ $\leq\left\|\eta\right\|_{L^{\infty}}^{2}+\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+2\mathbb{E}\left[\int_{t}^{T}\int_{\tilde{\Omega}}|g(s,z_{s},\tilde{z}_{s})-g(s,0,0)|d\tilde{\mathbb{P}}ds\Big|\mathcal{F}_{t}\right]^{2}.$ (6.34)
Together with the $(\rm H_{q}^{\prime})$} condition and Jensen’s inequality, we have
$\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|\leq L_{z}\big(|z_{s}|+\mathbb{E}|z_{s}|\big)+2K_{z}\big(|z_{s}|^{2}+\mathbb{E}|z_{s}|^{2}\big).$
Thus,
$\mathbb{E}\left[\int_{t}^{T}\left|\int_{\tilde{\Omega}}g(s,z_{s},\tilde{z}_{s})-g(s,0,0)d\tilde{\mathbb{P}}\right|ds\mid\mathcal{F}_{t}\right]$ (6.35)
$\leq\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right].$
Combining (6.34) and (6.35) gives
$\frac{1}{2}\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+2\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]\right]^{2}$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+4\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left(\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]^{2}+\mathbb{E}\left[\int_{t}^{T}L_{z}|z_{s}|+2K_{z}|z_{s}|^{2}ds\right]^{2}\right)$
$\leq\left\|\eta\right\|_{L^{\infty}}^{2}+16\left(L_{z}^{2}T\left\|z\right\|_{BMO}^{2}+4K_{z}^{2}\left\|z\right\|_{BMO}^{4}\right).$
The last inequality uses Jensen’s inequality:
$\left(\int_{t}^{T}L_{z}|z_{s}|\,ds\right)^{2}\leq(T-t)\left(\int_{t}^{T}L_{z}^{2}|z_{s}|^{2}\,ds\right).$
Now, using $a^{2}+b^{2}\leq(|a|+|b|)^{2}$, it follows that
$\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}$ $\leq 2\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+4\left\|Z\right\|_{BMO}^{2}$
$\leq 4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})+\beta^{2}(\left\|u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|z\right\|_{BMO}^{2})^{2},$
where $\alpha:=8L_{z}\sqrt{T}$, $\beta:=16K_{z}$. To obtain a uniform bound, set $R>0$ such that
$4\left\|\eta\right\|_{L^{\infty}}^{2}+\alpha^{2}R^{2}+\beta^{2}R^{4}\leq R^{2},$
which holds provided $\alpha^{2}-1<0$ (that is, $T<(64L_{z}^{2})^{-1}$), and we require
$\left\|\eta\right\|_{L^{\infty}}\leq\frac{1-\alpha^{2}}{4\beta}.$ (6.36)
In this case, we may take
$R^{2}=\frac{8}{(1-\alpha^{2})}\left\|\eta\right\|_{L^{\infty}}^{2}.$ (6.37)
The ball
$\mathcal{B}_{R}:=\left\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{BMO}:\left\|U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|Z\right\|_{BMO}^{2}\leq R^{2}\right\}$
is then such that $\Gamma(\mathcal{B}_{R})\subset\mathcal{B}_{R}$.
Step 2. Next, we prove that $\Gamma$ is a contraction on $\mathcal{B}_{R}$. Let $(u^{i},z^{i})\in\mathcal{B}_{R}$ ($i=1,2$), with $(U^{i},Z^{i})=\Gamma(u^{i},z^{i})$, and set
$\Delta u=u^{1}-u^{2},\quad\Delta z=z^{1}-z^{2},\quad\Delta U=U^{1}-U^{2},\quad\Delta Z=Z^{1}-Z^{2}.$
A similar computation as above yields
$\frac{1}{2}\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}$
$\leq 2\,\operatorname*{ess\,sup}_{(t,\omega)\in[0,T]\times\Omega}\left[\mathbb{E}\left(\int_{t}^{T}|g(s,z^{1}_{s},\tilde{z}^{1}_{s})-g(s,z^{2}_{s},\tilde{z}^{2}_{s})|ds\mid\mathcal{F}_{t}\right)\right]^{2}$
$\leq 4L_{z}^{2}T\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}|\Delta z_{s}|^{2}+\mathbb{E}|\Delta z_{s}|^{2}ds\mid\mathcal{F}_{t}\right]$
$\quad+16K_{z}^{2}\,\operatorname*{ess\,sup}_{(t,\omega)}\mathbb{E}\left[\int_{t}^{T}(|z^{1}_{s}|^{2}+|z^{2}_{s}|^{2}+\mathbb{E}|z^{1}_{s}|^{2}+\mathbb{E}|z^{2}_{s}|^{2})ds\mid\mathcal{F}_{t}\right]\mathbb{E}\left[\int_{t}^{T}(|\Delta z_{s}|+\mathbb{E}|\Delta z_{s}|)^{2}ds\mid\mathcal{F}_{t}\right]$
$\leq\left(8L_{z}^{2}T+32K_{z}^{2}(\left\|z^{1}\right\|_{BMO}^{2}+\left\|z^{2}\right\|_{BMO}^{2})\right)\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
Since $\left\|z^{i}\right\|_{BMO}^{2}\leq R^{2}$ for $i=1,2$, we obtain
$\left\|\Delta U\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta Z\right\|_{BMO}^{2}\leq(16L_{z}^{2}T+128K_{z}^{2}R^{2})\left(\left\|\Delta u\right\|_{\mathbb{S}^{\infty}}^{2}+\left\|\Delta z\right\|_{BMO}^{2}\right).$
Now, recall from the previous step that $R^{2}=\frac{8}{(1-\alpha^{2})}\left\|\eta\right\|_{L^{\infty}}^{2}$. In order for $R^{2}$ to be positive, we require $T<(64L_{z}^{2})^{-1}$, which is more restrictive than the condition here. Furthermore, if
$\left\|\eta\right\|_{L^{\infty}}^{2}<\frac{(4-\alpha^{2})(1-\alpha^{2})}{16\beta^{2}},$ (6.38)
then it follows that
$16L_{z}^{2}T+128K_{z}^{2}R^{2}<1,$
and thus the mapping $\Gamma$ is a contraction and admits a unique fixed point in $\mathcal{B}_{R}$.
Step 3. Finally, together with the local condition $T<(64L_{z}^{2})^{-1}$, the smallness condition (6.36) is in fact more stringent than the smallness condition (6.38). Consequently, if we define $\rho>0$ by
$\rho^{2}:=\frac{1-64L_{z}^{2}T}{65K_{z}},$
then, provided that $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, the general BSDE admits a unique adapted solution $(U,Z)\in\mathcal{B}_{\rho}$ on this local time interval, under the imposed smallness conditions on both the terminal value and the time horizon. This completes the proof of local existence and uniqueness. ∎
Proposition 6.7.
Let $T>0$, and assume that $g$ satisfies $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. Suppose there exists $\rho>0$ such that
$\left\|\eta\right\|_{L^{\infty}}\leq\rho,$
where
$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
Then the BSDE (6.33) admits a unique solution $(U,Z)\in\mathcal{B}_{\rho}$ on $[0,T]$, where
$\mathcal{B}_{\rho}:=\Bigl\{(U,Z)\in\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}:\ \|U\|_{\mathbb{S}^{\infty}}^{2}+\|Z\|_{BMO}^{2}\leq\rho^{2}\Bigr\}.$
Proof.
Let $\delta:=(65L_{z}^{2})^{-1}<(64L_{z}^{2})^{-1}$, and divide $[0,T]$ into $N$ subintervals of length at most $\delta$: $0=t_{0}<t_{1}<\cdots<t_{N}=T$, where $t_{k+1}-t_{k}\leq\delta$ for all $k$. By construction, $\rho^{2}=\frac{1-64L_{z}^{2}\delta}{65K_{z}}=\frac{1}{65^{2}K_{z}}$. We construct the solution recursively, starting from the last interval $t_{N-1},t_{N}]$. On this interval, the terminal value is $\eta$, and the smallness condition is satisfied by assumption. Thus, by applying the local existence and uniqueness lemma, we obtain a unique solution $(U,Z)$ on $t_{N-1},t_{N}]$ with a priori bound in $\mathcal{B}_{\rho}$. Then we proceed recursively: at each interval $t_{k-1},t_{k}]$, we solve the BSDE using $U_{t_{k}}$ as the terminal value. The a priori bound $\|U\|_{\mathbb{S}^{\infty}}\leq\rho$ ensures that the smallness condition $\|U_{t_{k}}\|_{L^{\infty}}\leq\rho$ continues to hold at each step. Finally, By concatenating these local solutions, we obtain a unique adapted solution $(U,Z)$ on $[0,T]$. ∎
We remark that, given the unique solution $(U,Z)\in\mathcal{B}_{\rho}$ constructed above for the auxiliary BSDE 6.33, the solution to the original BSDE 6.32 is obtained by the explicit transformation
$Y_{t}=U_{t}+g(t,0,0)+\mathbb{E}[\xi],\qquad 0\leq t\leq T.$
Furthermore, we emphasize that the smallness condition $\left\|\eta\right\|_{L^{\infty}}\leq\rho$, with $\eta:=\xi-\mathbb{E}[\xi]$, requires the terminal condition $\xi$ to be close to its mean. This is a slightly more general requirement than simply demanding that $\xi$ itself be small, as it allows for arbitrary (but fixed) expectations and only controls the oscillation of $\xi$ around its mean.
Remark 6.8.
Proposition 6.7 was established for drivers of the form $\int_{\tilde{\Omega}}g(s,Z_{s},\tilde{Z}_{s})\,d\tilde{\mathbb{P}}$, which naturally extends to the case where the driver depends on the expectation $\mathbb{E}[Z]$ instead. In particular, the existence and uniqueness result applies to BSDEs of the type
$Y_{t}=\xi+\int_{t}^{T}g(s,Z_{s},\mathbb{E}[Z_{s}])\,ds-\int_{t}^{T}Z_{s}\!\cdot dW_{s},\quad t\in[0,T],$
which will be useful for analyzing the driver (6.31) we picked earlier in the context of utility maximization in this section.
While the above analysis focuses on drivers with $Z$ and mean-field dependence, it is very easy to extend existence and uniqueness results to more general settings where the driver $g$ may also depend on the process $Y$ and its expectation $\mathbb{E}[Y]$. In the present section, however, the application of interest involves drivers depending solely on $Z$ and its expectation, so the above results apply directly. We now present the main result of this section.
Theorem 6.9.
Suppose the liability $F$, the terminal condition of BSDE (6.30) satisfies the condition such that $\left\|F-\mathbb{E}[F]\right\|_{L^{\infty}}\leq\rho$ where
$\rho^{2}=\frac{1}{65^{2}K_{z}}.$
Let $(Y,Z)$ denote the unique solution to BSDE (6.30) with driver $f$ (6.31). Then the value function $V^{(p)}(t,x)$ associated with the utility maximization problem over the admissible set $\mathcal{A}$ and constraint set $C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)$ is given by
$V^{(p)}(0,x)=-\exp(-\alpha(x-Y_{0})),$
and there exists an optimal trading strategy $p^{*}\in\mathcal{A}$ such that, for all $t\in[0,T]$ and $\omega\in\Omega$,
$p^{*}_{t}(\omega)\in\Pi_{C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)}\left(Z_{t}+\frac{\theta_{t}}{\alpha}\right).$
Proof.
First we will show the driver $f$ (6.31) satisfies the condition $(\mathrm{H}_{0})$ and $(\mathrm{H}_{q}^{\prime})$. The verification of the first condition is straightforward due to the boundedness of $\theta$. To verify $(\mathrm{H}_{q}^{\prime})$, we apply Lemma 6.3. For any $\omega\in\Omega$ and $t\in[0,T]$, since $Z_{t}(\omega)\in\mathbb{R}^{d}$, we have
$\text{dist}^{2}\left(Z_{t}+\frac{\theta_{t}}{\alpha},\,C_{t}\left(\omega,\mathbb{E}[Z_{t}]\right)\right)$ $\leq 2|Z_{t}|^{2}+2\left(\frac{|\theta_{t}|}{\alpha}+K_{1}\mathbb{E}|Z_{t}|\right)^{2}$
$\leq 2\frac{|\theta_{t}|^{2}}{\alpha^{2}}+\frac{K_{1}|\theta_{t}|}{\alpha}+2|Z_{t}|^{2}+2\left(K_{1}^{2}+\frac{K_{1}|\theta_{t}|}{\alpha}\right)\mathbb{E}|Z_{t}|^{2}.$
Now, for $t\in[0,T]$ and $Z_{t}^{1}(\omega),Z_{t}^{2}(\omega)\in\mathbb{R}^{d}$, we observe
$f(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}])-f(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}])$
$\qquad=-\frac{\alpha}{2}\left[\text{dist}^{2}\left(Z_{t}^{1}+\frac{\theta_{t}}{\alpha},\,C_{t}(\omega,\mathbb{E}[Z_{t}^{1}])\right)-\text{dist}^{2}\left(Z_{t}^{2}+\frac{\theta_{t}}{\alpha},\,C_{t}(\omega,\mathbb{E}[Z_{t}^{2}])\right)\right]+(Z_{t}^{1}-Z_{t}^{2})\theta_{t}.$
By the Lipschitz property of the distance function from a closed set, it follows that, for all $t\in[0,T]$
$\left|f\left(t,Z_{t}^{1},\mathbb{E}[Z_{t}^{1}]\right)-f\left(t,Z_{t}^{2},\mathbb{E}[Z_{t}^{2}]\right)\right|$
$\quad\leq\frac{\alpha}{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+2\frac{|\theta_{t}|}{\alpha}+K_{1}\big(\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\big)\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+K_{1}\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right)+|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|$
$\quad\leq 2|\theta_{t}|\,|Z_{t}^{1}-Z_{t}^{2}|+\frac{\alpha}{2}(1\vee K_{1})^{2}\left(|Z_{t}^{1}|+|Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}|+\mathbb{E}|Z_{t}^{2}|\right)\left(|Z_{t}^{1}-Z_{t}^{2}|+\mathbb{E}|Z_{t}^{1}-Z_{t}^{2}|\right),\quad\text{a.s.}$
Together with the boundedness of $\theta$, we obtain a sufficient condition for $(\rm H_{q}^{\prime})$}. This ensures the existence and uniqueness of a solution $(Y,Z)$ to BSDE (6.30) by Proposition 6.7. It remains to show that for any $p\in\mathcal{A}$, the process
$V^{(p)}(t,x):=-\exp\left(-\alpha(X_{t}^{p}-Y_{t})\right),\quad t\in[0,T],$
is a supermartingale, and that it becomes a martingale when $p=p^{*}$ as defined above. This verification follows directly from arguments in Theorem 7 of hu2005utility, and we omit the details here. ∎
7 Conclusion
In this paper we have developed a general framework for mean-field BSDEs whose generators have quadratic growth in the control and its law. Working under the separately quadratic structure (Hq) with bounded terminal condition, we proved global existence and uniqueness of solutions in the Banach space $\mathbb{S}^{\infty}\times\mathbb{H}^{2}_{\mathrm{BMO}}$, without imposing any smallness condition. The key ingredients were a Malliavin calculus argument for Lipschitz mean-field BSDEs, which yields uniform bounds on the control process, together with BMO estimates and a global stability result that allows us to pass from Lipschitz to quadratic generators. We then extended the analysis to the fully coupled quadratic regime (Hq), in which the generator admit genuine cross-interaction terms. In this more general setting we established existence and uniqueness under a natural smallness condition on the centred terminal condition. As an application, we showed how the fully coupled theory can be used to formulate and solve a mean-field exponential utility maximization problem, thereby generalizing the classical framework of hu2005utility.